OpenStudy (anonymous):

Trevor's dad decided to get him a cell phone for his birthday. Cell Plus has a plan that costs $24.00 per month plus an additional $0.06 per minute. Cell Best has a plan that costs $30.00 per month plus an additional $0.05 per minute. How many minutes can Trevor talk and both cell plans cost the same amount? 600 60 300 120

OpenStudy (anonymous):

24*m+.06min=cell plus 30*m +. 05min=cell best

OpenStudy (anonymous):

if you make the ms 1 and the mins xs then you have a y=mx+b equation y=.06x+24 y=.05x+30 now just solve you can view it as plugin in the y into the other. .06x+24=.05x+30

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

Trevor's dad decided to get him a cell phone for his birthday. ----- Cell Plus has a plan that costs $24.00 per month plus an additional $0.06 per minute. that can be interpreted as 24*month +.06 minutes=cellplus's prize --- Cell Best has a plan that costs $30.00 per month plus an additional $0.05 per minute. that can be interpreted as 30*month+.05*minute= cellbest's prize --- How many minutes can Trevor talk and both cell plans cost the same amount? first we put both math interpratations one on top of the other. 24*month +.06* minutes=cellplus's prize 30*month+.05*minute= cellbest's prize We assume that in 1 month how many minutes can Trevor talk and both cell plans cost the same amount? we plug in 1 where the months are at 24*1 +.06* minutes=cellplus's prize 30*1 +.05*minute= cellbest's prize SInce they ask for when are both prices the same. THe math interpretation is cellplus's prize=cellbest's prize We plug in therefore 24*1 +.06* minutes=30*1 +.05*minute But in the other problem we just uses x's in place of minutes :) and y's in place of prizes. You just solve for minutes.

OpenStudy (anonymous):

so its C?

OpenStudy (anonymous):

i get 600

OpenStudy (anonymous):

.06x+24=.05x+30 .06x=.05x+6 subtract 24 both sides .01x=6 subtract .05 both sides :) x=6/.01 which is 600 when you divide.

OpenStudy (anonymous):

thank you!